Please show work
Letf(x)=5x2 –6 and g(x)=x–2.
(a) Find the composite function ( f o g)(x) and simplify. Show work.
(b)Find (f og)(1)

Answer by Cromlix(4381)

Hi there,
f(x) = 5x^2 g(x) = x - 2
(f o g)(x) = 5(x - 2)^2
(f o g)(x) = 5(x^2 - 4x + 4)
(f o g)(x) = 5x^2 - 20x + 20
b) (f o g)(1) = 5(1)^2 - 20(1) + 20
(f o g)(1) = 5 - 20 + 20
(f o g)(1) = 5
Hope this helps :-)